### v1

# HAAP: Vision-context Hierarchical Attention Autoregressive with Adaptive Permutation for Scene Text Recognition

Abstract
Internal Language Model (LM)-based methods use permutation language modeling (PLM) to solve the error correction caused by conditional independence in external LM-based methods. However, random permutations of human interference cause fit oscillations in the model training, and Iterative Refinement (IR) operation to improve multimodal information decoupling also introduces additional overhead. To address these issues, this paper proposes the Hierarchical Attention autoregressive Model with Adaptive Permutation (HAAP) to enhance the location-context-image interaction capability, improving autoregressive generalization with internal LM. First, we propose Implicit Permutation Neurons (IPN) to generate adaptive attention masks to dynamically exploit token dependencies. The adaptive masks increase the diversity of training data and prevent model dependency on a specific order. It reduces the training overhead of PLM while avoiding training fit oscillations. Second, we develop Cross-modal Hierarchical Attention mechanism (CHA) to couple context and image features. This processing establishes rich positional semantic dependencies between context and image while avoiding IR. Extensive experimental results show the proposed HAAP achieves state-of-the-art (SOTA) performance in terms of accuracy, complexity, and latency on several datasets.
Index Terms:
Language models (LM), Autoregressive generalization, Multimodal information.
I Introduction
Scene Text Recognition (STR) aims to transcribe the text in an image into a computer-readable text format i.e., to recognize the localized text regions. Scene text with rich information plays a vital role in a range of applications such as visual quizzing, autonomous driving, image retrieval, augmented reality, retail, education, and visually impaired devices [1, 2, 3, 4, 5].
Figure 1: Illustration of IPN. The solid line represents the process of mask generation i.e. non-linear weighted mapping of left-to-right permutations. The dashed line represents the interpretation of the mask generation: the visual information guides the adaptive mask to learn the inter-correlation of the contextual positions.
In contrast to Optical Character Recognition (OCR) in documents with more homogeneous text attributes, STR has to deal with irregular text, such as rotation, bending, blurring, or occlusion. In addition, noise and distortion issues make STR challenging. STR is primarily a visual task, involving two different modalities: image and text. Early work still tends to rely on a backbone pre-trained on unimodal data. Learning from data [6, 7, 8, 9, 10, 11, 12, 13] is achieved by combining visual features with semantic priors from word representation models [14] or dictionaries [15] as well as sequence modeling [16]. However, there are inherent limitations in models based on visual information. Visual features alone are not sufficient in cases where the text part is unreadable, e.g., occlusion. Additionally, STR is a One-shot task that has data independence. And, textual semantics can assist visual models in recognizing heterogeneous data.
Zhao et al. verified that language modeling (LM) can effectively focus on text and accurately determine the meaning of words [13]. LM demonstrates the special ability to perceive and understand text in natural images. For that, the combination of LM and visual models for sequence modeling is the current mainstream [8, 9, 10, 11, 12, 13]. It is divided into two categories: external LM-based and internal LM-based. External LM-based schemes essentially accomplish STR by introducing external LMs to assist in correcting visual model reasoning [8, 9, 10, 11]. External LM causes error correction own to the conditional independence of the input. The condition assumes that each pixel or character is independent, regardless of their interactions. However, the reality is that the order and spatial position of the characters about each other affect their meaning and grammatical structure. Algorithms that only consider each character by itself and ignore interactions can lead to prediction errors, especially when the input text has a complex structure and arrangement. Therefore, Permutation Language Modeling (PLM) is employed to improve the generalization of autoregressive (AR) modeling by parsing sequence meaning and syntactic structure [12, 13]. PLM models can be viewed as a collection of AR models with shared architecture and weights. Token dependencies are specified dynamically through the use of an attention mask. Although randomized permutation increases the diversity of data, this strategy is accompanied by training fit oscillations since the random subset has the uncertainty of human interference. We also find that random permutation leads to unstable convergence of the model during training. In addition, if the model only performs well for specific alignments, it leads to a degradation of the generalization performance. Next, recent approaches have used Transformer encoder-decoder based for cross-modal visual and semantic information fusion [8, 9, 10, 11, 12, 13]. However, They need to optimize the predictions through additional Iterative Refinement (IR). Essentially, internal LM is introduced in the decoder can eliminate IR, which is only an additional decoding because the visual and LM are jointly parsed. Considering that noisy inputs still have errors from the output of the visual model, IR is proposed to refine predictions from visual and language cues in external LM-based scenarios. Therefore, recent internal LM-based schemes do not propose an efficient way to couple cross-modal information in the decoder but rely on additional IR.
Figure 2: The basic flow of STR. (a) Visual feature coding and decoding (b) Visual-contextual representation based on external LM. (c) Visual-contextual representation based on internal LM. (d) Internal LM-based visual-contextual representation without Iterative Refinement (IR) (Ours).
This paper proposes the Hierarchical Attention AR Model with Adaptive Permutation (HAAP) to address these issues. The motivation is depicted. First, the use of an adaptive permutation instead of a random permutation scheme increases the diversity of the training data and prevents dependence on a specific order. It reduces the training overhead of the PLM while avoiding training fitting oscillations (as shown in Fig. 1). Second, Multi-head Attention (MHA) is introduced to hierarchically couple cross-modal information. Hierarchical processing improves the adaptation of visual features to semantics and thus enhances end-to-end joint parsing without additional IR (as shown in Fig. 2).
First, inspired by traditional neurons and self-attentive mechanisms, this paper proposes an Implicit Permutation Neurons (IPN) to adapt the ordering between characters in the internal LM. The ordering and spatial location of the characters are considered rather than independently. IPN assigns a set of queries to the original left-to-right permutation to map it linearly into the high-dimensional space. Immediately, two sets of weight matrices are utilized for query weighting and inverse mapping. The process is still able to capture nonlinear relationships in the data with learnable weights. Second, this paper proposes Cross-modal Hierarchical Attention mechanism (CHA) to couple cross-modal information. Contexts embedded in the high dimensional space establish semantic dependencies through self-attention. The context interacts with the image features for the first time to context-image feature alignment. Next, the encoded location is matched with the context and then coupled with the image features again by mutual attention. Hierarchical feature processing drives the interactions of each item in the sequence while capturing the dependencies between locations using an image-guided context-assisted model. In addition, the location-context-image interaction is improved with location inter-correlation guided by adaptive mask and the uniform gradient flow. Therefore, the model avoids additional overhead by eliminating IR and achieves end-to-end parsing.
The contributions of this paper can be summarized as follows:
  • This paper proposes an internal LM-based HAAP to unify end-to-end coupling of location coding, image features, and context to improve the generalization ability of the model in various STR scenes.
  • We propose a novel neuronal structure i.e. IPN to capture the dependencies between characters via adaptive masks efficiently. The model automatically learns the optimal alignment form by capturing the nonlinear relationships during the learning of weights without manual design.
  • We propose hierarchical feature processing CHA to motivate the interaction of each item in the sequence, together with adaptive masks to improve the location-context-image interaction.
  • HAAP achieves state-of-the-art (SOTA) results on the STR benchmark for all character sets as well as on larger, more difficult real datasets that include occluded and arbitrarily oriented text (as shown in Table. II). Besides, HAAP also shows the cost-quality trade-offs in terms of parameters, FLOPs, and runtime usage (as shown in Table. IV).
The rest of the paper is structured as follows: section II discusses related work. Section III describes the architecture of the proposed algorithm in detail. Section IV describes the experiments and analysis and the summary in Section V.
II Related work
In this section, we review current deep learning-based STR methods. Recent studies are mainly categorized into two directions, context-independent and context-relevant, based on whether textual semantics are considered or not. We summarize and discuss the research in these two directions separately.
II-A Context-independent STR
Considering that STR is essentially a visual task, the researchers intuitively represent the inputs by constructing a visual model based on deep learning. This is the context-independent scheme i.e., category inference by summarizing visual features. Further, context-independent STR can be categorized into parallel-based reasoning and sequential reasoning. The parallel-based scheme only uses the visual features for prediction without considering the relationship between characters. Its main application is Fully Convolutional Network (FCN) [17] to segment characters at the pixel level. Liao et al [18] recognized characters by grouping segmented pixels into text regions. Wan et al [19] proposed an additional sequential segmentation map that records characters in the correct order. Multi-category classification is unable to accurately construct complete phrases because the output characters are conditionally uncorrelated with each other.
Immediately, the scheme of sequence inference using Recurrent Convolutional Networks (RNN) was proposed to capture the correlation of characters [20]. The most typical is the Connectionist Temporal Classification (CTC) [21] based scheme. The RNN models the sequential sequence of features extracted by a convolutional neural network (CNN) and is trained end-to-end using CTC loss [22]. The RNN receives the current input and the previous hidden state at each time step and outputs the current hidden state. This sequential modeling structure allows the RNN to naturally adapt to the characteristics of sequential data, to parse the syntactic and semantic structures in the text. RNN is also used in a series of Attention-based [23] schemes. Several attempts have explored the enhancement of image encoding and decoding by employing different attentional mechanisms [24, 25, 26]. Character classification by aligning features and character positions, or converting STR into a multi-instance learning problem [7, 27]. Sequence modeling was given a new lease of life with the introduction of transformer [16, 28]. Transformer was then proposed as an alternative to RNN for sequence modeling [29, 30, 31, 32, 33, 34, 35]. However, context-independent methods that rely only on image features for prediction cannot address recognition in low-quality images, especially being less robust to corruption such as occlusion, distortion, or incomplete characters. This limitation motivates the use of language semantics to make recognition models more robust.
II-B Context-relevant STR
Context-dependent STR is a typical cross-modal fusion scenario since the internal interactions between vision and language are constructed. Considering that textual semantics can assist visual models in processing heterogeneous data, the community has recently appeared with a large number of researches to assist recognition by constructing LMs [14, 8, 9, 12, 10, 36, 13, 11]. VisionLAN [8] introduced character masking to guide the visual model to refer to language information in the visual context. MATRN [10] referenced context to enhance visual and semantic features based on spatial coding. STRT [11] designed an iterative text Transformer to predict the probability distribution in a sequence of characters. It is shown that robust models can be constructed by introducing external LMs for explicit modeling. Considering that character sequences are usually modeled in a left-to-right manner, the researcher develops the integrated model to capture twice the amount of information through the bi-directional LM [7, 37, 6, 38]. SRN [7] combined the features of two unidirectional Transformers for prediction, which resulted in twice as expensive both computationally and parametrically. Therefore, ABINet [9] used a novel bidirectional completion network based on bidirectional feature representation. However, the conditional independence of external LM on image features limits its performance and makes error corrections. Subsequently, the internal LM was re-proposed for implicit language modeling to improve visual-contextual interaction. PARSeq [12] used PLM rather than standard AR modeling to learn internal LMs. It supports flexible decoding by using a parameterization that decouples the decoding location from the input context. CLIP4STR [13] introduced an encoder-decoder to process image and text information. The encoder is inherited from CLIP [39], while the decoder uses Transformer.
Internal LR-based schemes mitigate the problem of noisy inputs as well as increase the diversity of the data by dynamically specifying the sequence order through the use of an attention mask. However, the strategy of using randomized permutations introduces training fitting oscillations due to the uncertainty of randomized permutations with human interference. Furthermore, most of these context-dependent STR models use Transformer-based encoder-decoders to implement with IR. However, internal LM-based schemes rely on IR to address noisy inputs in external LM-based models. Thus, the decoding strategy of LM model-based schemes at this stage is inefficient and introduces additional computational effort. This paper proposes to use an adaptive sequence ordering scheme to increase the diversity of the training data and prevent the dependence of the model on a specific order. To further enhance the efficiency of cross-modal information coupling at the decoder side, MHA is developed to Hierarchical coupling information, improving the semantic adaptation of visual features. Ultimately, end-to-end joint parsing is enhanced to avoid additional IR.
Figure 3: The pipeline of HAAP. (a) Image and text inputs are represented as a series of patches and semantic tokens. (b) Visual information guides the IPN to assign bidirectional and adaptive masks to the context. (c) MHA is used to perform hierarchical image-context interaction and decoding.
III Methodology
In this section, we present in detail the principles and framework of our proposed Hierarchical Attention AR Model with Adaptive Permutation (HAAP), including Implicit Permutation Neurons (IPN) and Cross-modal Hierarchical Attention mechanism (CHA).
III-A Overview
HAAP follows an encoder-decoder architecture as shown in Fig. 3. In the encoding phase, the image and text inputs are represented as a series of patches and semantic tokens. Subsequently, MHA is used for hierarchical image-context interaction and decoding. Specifically, first, the Transformer encoder [28] is utilized to establish the internal feature associations of the image patches. Second, the IPN assigns bi-directional and adaptive permutation masks to the textual context for adjusting sequence order. The contexts embedded into the high-dimensional features are first established by self-attention to establish internal associations. Then MHA is used for context and visual information interaction and position query association. Third, the associated information further interacts with the visual information by utilizing mutual attention. Next, the position query is linearly decoded into contextual output.
Figure 4: Illustration of a ViT layer from Dosovitskiy et al . [28]. $LN$ pertains to layer normalization. $MLP$ represents Multilayer Perceptron.
III-B Encoder
Vision encoder Considering that ViT [28] divides images into patches and uses full connectivity to learn global relationships can better extract global and complex visual information than traditional CNNs [40], HAAP follows the original ViT structure to build the visual encoder, which consists of $L=12$ layers of Transformer encoder. All layers share the same architecture as shown in Fig. 4. The output of the last encoder is subjected to a Layer Normalization (LN). Formally, first, the image of size $H\times W\times C$ (height $H$, width $W$, channels $C$) is reshaped into a series of flat 2D patches $x_{p}\in\mathbb{R}^{N\times(P^{2}-C)}$, where $(P,P)$ is the resolution of each image block and $N=HW/P^{2}$ is the number of blocks generated. The patch $x_{p}$ is flattened and mapped to $D$ dimensions using a trainable linear projection $E\in\mathbb{R}^{({P^{2}}C)\times D}$. The position embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ is added to the projected output $z$, as:
$z=[x^{0}_{p}E;x^{1}_{p}E;x^{2}_{p}E;\cdots;x^{N}_{p}E;]+E_{pos},$ (1)
where $x^{i}_{p}$ represents the $i$-th $x_{p}$. $z$ performs feature extraction in alternating MHA and Mult-Layer Perceptron (MLP). A normalized $LN$ is applied initially in each block to mitigate internal covariate bias during training. Additionally, residual concatenation is applied twice in the block to aid gradient propagation in the depth model, as:
$z^{\prime}_{l}=MHA(LN(z_{l-1}))+z_{l-1},l\in[1,L]$ (2)
$z_{l}=MLP(LN(z^{\prime}_{l}))+z^{\prime}_{l},l\in[1,L]$ (3)
where $MLP$ contains two nonlinear layers with GELU activation functions. The $z^{\prime}_{l}$ and $z_{l}$ are the visually encoded information. MHA is the extension of scaled dot-product attention to multiple representation subspaces or heads. In scaled dot-product attention, the similarity scores are computed using dot-product computation of the similarity between two $d_{k}$-dimensional vectors query $q$, and key $k$. Attention scores are transformed by the $d_{v}$-dimensional vector values $v$. The scaled dot product attention is defined as:
$Attn(q,k,v)=softmax(\frac{qk^{T}}{\sqrt{d_{k}}})v,$ (4)
further,
$MHA(q,k,v)=Concat(head_{1},\cdots head_{h})W^{o},$ (5)
where
$head_{i}=Attn(qW^{q}_{i},kW^{k}_{i},vW^{v}_{i}),$ (6)
where the computational cost of MHA is practically constant due to the dimension of the vector $d_{head}=d_{model}/h$, where $h$ is the number of heads. The $W^{q},W^{k},W^{v}\in\mathbb{R}^{d_{model}\times d_{head}}$ are obtained by connecting the heads and multiplying with the output projection matrix $W^{o}\in\mathbb{R}^{d_{model}\times d_{head}}$.
Context encoder The context uses a Tokenizer [12] for lowercase byte pair encoding BPE. The start, padding, and end of the text sequence are padded with [SOS], [PAD] and [EOS] tokens, respectively.
III-C Implicit Permutation Neurons (IPN)
Considering that STR traditionally relies on left-to-right $P_{left2right}$ or right-to-left $P_{right2left}$ unidirectional sequence modeling, we propose IPN to improve the correlation between inputs and outputs through multidirectional sequence modeling. Essentially, multidirectional sequence modeling considers internal correlations while considering bi-directionality to avoid the effects of word syntactic structures or external disturbances such as occlusions and distortions. For example, to predict the ”e” in the word ”clearance”, we consider ”cl_” and ”ecnara_”. ” in the context of ”cl_ar”. IPN provides an attention mask during the attention operation to generate dependencies between the input context and the output without actually replacing the text labels. Table. I illustrates three examples of the mask.
TABLE I: Illustration of Masks from IPN. $P_{left2right}$ and $P_{right2left}$ represent the left-to-right and right-to-left permutation.[B] And [E] represent the beginning and end of the sequence. $y1,y2,y3$ represent the three positions of this sequence. 1 indicates that the output has a conditional dependency on the input. 0 indicates that there is no information flowing from input to output.
$P_{left2right}$ Mask
[B] y1 y2 y3
y1 1 0 0 0
y2 1 1 0 0
y3 1 1 1 0
[E] 1 1 1 1
$P_{right2left}$ Mask
[B] y1 y2 y3
y1 1 0 1 1
y2 1 0 0 1
y3 1 0 0 0
[E] 1 1 1 1
Attention mask of IPN
[B] y1 y2 y3
y1 1 0 1 1
y2 1 1 0 0
y3 1 1 1 0
[E] 1 1 1 1
Formally, first, given the $T$-length text labels $y=[y^{1},y^{2},\cdots,y^{T}]$, decompose the likelihood $[1,2,\cdots,T]$ according to the canonical order using the chain rule to build a left-to-right permutation of $P_{left2right}$ to obtain the model inference as:
$\mathrm{log}p(y|x)=\sum_{t=1}^{T}{\mathrm{log}p_{\theta}(y_{t}|y_{<t},x)},$ (7)
where $\theta$ represents the model parameter and $x$ is the input image. Second, to learn and extract complex structural properties from the data, $P_{query}$ is assigned to $T$ positions in the $P_{left2right}$ sequence and projected to a high-dimensional space, as:
$M_{t}=P_{left2right}^{\intercal}\cdot P_{query},$ (8)
where $M_{t}\in\mathbb{R}_{T}^{T}$ represents the intermediate matrix.
Third, the model learns the optimal representation without manually designing features via a learnable weight matrix $P_{weight}$, as:
$M_{w}=M_{t}\cdot P_{weight},$ (9)
where $M_{w}$ is the weighted transformation matrix. This process helps the model to parse the representation of the structure of the input data. Even though there is no explicit nonlinear activation function throughout the process, the model is still able to capture the nonlinear relationships in the data because of the learning of the weights. Fourth, the transformation matrix is inversely mapped to the sequence of scores $P_{score}$ via $P_{value}$, as:
$P_{score}=M_{w}\cdot P_{value}^{\intercal}.$ (10)
Next, $P_{score}$ is normalized from largest to smallest to generate the actual ranking order. Immediately, we invert the left-to-right permutations $[1,2,.,T]$ and invert the adaptive permutation pairs to obtain twice the amount of information.
III-D Cross-modal Hierarchical Attention mechanism (CHA)
The standard Transformer parametric AR model is inefficient for decoding in multidirectional sequence modeling since the feature distribution predicted by the hidden state is independent of the target location [41]. Let us assume that we parameterize the next token distribution as:
$p_{\theta}(x_{rt}=x|x_{r<t})=\frac{exp(e(x)^{\intercal}h_{\theta}(x_{r<t}))}{\sum{x^{\prime}}exp(e(x\prime)^{\intercal}h_{\theta}(x_{r<t}))},$ (11)
where $h_{\theta}(x_{r<t})$ denotes the hidden representation of $x_{r<t}$ produced by the shared transformer network after proper masking. We use $rt$ and $r<t$ to denote the $t$-th and the first $t-1$ elements of the permutation set $r$. Since the same model parameter $\theta$ is shared between all decomposition sequences during training, $rt$ is expected to see that every possible element in the sequence is not equal to $x_{t}$. Note that the representation $h_{\theta}(x_{z<t})$ does not depend on the position it will predict, i.e. the value of $rt$. Two different target positions share the same model predictions. However, the ground truth distributions for the two positions are certainly different.
To avoid this problem, we perform context and location encoding based on a two-stream attention mechanism to provide complete contextual information. We use two sets of hidden representations i.e., from content and query representations. Meanwhile, considering that context and location need to interact with visual features cross-modally for alignment, we adopt a hierarchical coupling approach, i.e., CHA, to consider the context-location-image association. In the first stage, context-visual attention is established after self-attentive encoding of sorted contexts, as:
$Attn_{cv}=p+dropout[MHA(Attn_{c},z)],$ (12)
where
$Attn_{c}=c+dropout[MHA(c,c,M)],$ (13)
where $c\in\mathbb{R}^{(T+1)\times d_{model}}$ is the contextual embedding with positional information and $M\in\mathbb{R}^{(T+1)\times(T+1)}$ is the attention mask. The total sequence length is increased to $T+1$ by using special separator markers.
Next, MHA is used to establish positional-contextual-visual attention, as:
$Attn_{f}=Attn_{p}+dropout[MHA(Attn_{p},z)],$ (14)
where
$Attn_{p}=p+dropout[MHA(p,Attn_{c}v,M)].$ (15)
The strategy of hierarchical feature processing prompts sequence interactions while utilizing image-guided context to assist the model in capturing dependencies between locations. The location-context-image interaction capability is boosted since adaptive permutation guides location interconnections. The robustness of the model is improved because the uniform gradient flow allows them to interact with each other during training. Finally, the output logits $y^{\prime}$ are obtained as:
$y^{\prime}=Linear(Attn_{f})\in R^{(T+1)\texttimes(S+1)},$ (16)
where $S$ is the size of the character set used for training. Additional characters are associated with the [E] tag (the end of the tagging sequence). The complete training loss for the optimization phase is the average of the $K$ cross-entropy loss $\mathcal{L}_{ce}$ with attention mask, as:
$\mathcal{L}=\frac{1}{k}\sum_{k=1}^{K}{\mathcal{L}_{ce}(y^{\prime}_{k},y)},$ (17)
where $y^{\prime}_{k}$ represents the $k$-th output logit and $y$ is the real label. $K=4$ is the number of permutations.
IV Experiments
IV-A Dataset
This experiment uses both synthetic training datasets MJSynth (MJ) [27] and SynthText (ST) [42] as well as real data for training including nine real-world datasets (COCO-Text (COCO) [43], RCTW17 [44], Uber-Text (Uber) [45], ArT [46], LSVT [47], MLT19 [48], ReCTS [49], TextOCR [50] and OpenVINO [51]). We use IIIT 5k-word (IIIT5k) [52], CUTE80 (CUTE) [53], Street View Text (SVT) [54], SVT-Perspective (SVTP) [55], ICDAR 2013 (IC13) [56] and ICDAR 2015 (IC15) [57], COCO, Uber, ArT for testing. The dataset contains challenging texts such as multi-orientation, curvature, perspective blur, low resolution, distortion, and occlusion.
Synthetic datasets MJ consists of 8.9 million photorealistic text images composed of artificial datasets through over 90k English dictionaries. The composition of MJ consists of three parts background, foreground, and optional shadow/border respectively. It uses 1400 different fonts. Additionally, the font word spacing, thickness, underlining, and other attributes of MJ are different. MJ also utilizes different background effects, border/shadow rendering, base shading, projection distortion, natural image blending, and noise. ST consists of 8 million word text images synthesized by blending text over natural images. It uses scene geometry, textures, and surface normals to naturally blend and distort text renderings on the surfaces of objects in the image. Similar to MJ, ST uses randomized fonts for its text. The text image is cropped from the natural image in which the synthesized text is embedded.
Real-world datasets The COCO dataset has a total of 73,127 training and 9.8k test images containing non-text, legible text, illegible text, and occluded text images. Each image contains at least one instance of legible text. RCTW contains 12,263 annotated images of the large-scale Chinese field dataset. The Uber dataset has a total of 127,920 training 60,000 test images collected from Bing Maps Streetside. It contains house numbers and vertical and rotated text on sign boards. The ArT dataset was created to recognize arbitrarily shaped text containing 32,028 training and 26,000 testing images of perspective, rotated, or curved arbitrarily shaped text. SVT is a large-scale street scene text dataset collected from streets in China totaling 41,439 text images of arbitrarily shaped natural scenes. MLT19 was created to recognize multilingual text. It has a total of 56,727 training images consisting of seven languages: Arabic, Latin, Chinese, Japanese, Korean, Bengali and Hindi. The ReCTS dataset contains 26,432 irregular texts arranged in various layouts or written in unique fonts. TextOCR and OpenVINO are large datasets with very diverse images containing 818,087 and 2,071,541 images respectively. They have complex scenes with multiple objects and text of different resolutions, orientations, and qualities.
The IIIT5k dataset is a collection of 5000 natural scene and digitally born text images crawled from Google image search containing 2,000 images for evaluation and 3,000 images for testing. It has low resolution, multiple font styles, light and dark transformations, and projection distortion. CUTE contains 288 cropped images for the curved text collection. The images were captured by digital cameras or collected from the Internet. SVT is a collection of street view text from Google Street View. It contains 257 images for evaluation and 647 images for testing. Further, the SVTP dataset is a more complex collection of images containing numerous perspective texts totaling 645 images for testing. IC13 and IC15 were created for the ICDAR Robust Reading Contest. IC13 contains 848 images for training and 1,015 images for evaluation. IC15 contains 4,468 images for training and 2,077 images for evaluation. Many of the samples contain perspective, and blurred text.
TABLE II: The comparison of word accuracy on six common benchmarks in terms of both context relevant and independent. “S” and “R” denote Synthetic and real data, respectively.
Method Train data IIIT5K(%) SVT(%) IC13(%) IC13(%) IC15(%) IC15(%) SVTP(%) CUTE(%)
Context-independent
CRNN(2016)[22] S 84.3 78.9 88.8 61.5 64.8 61.3
SRN(2020)[7] S 94.8 91.5 95.5 82.7 85.1 87.8
ViTSTR(2021)[29] S 88.4 87.7 93.2 92.4 78.5 72.6 81.8 81.3
TRBA(2021)[30] S 92.1 88.9 93.1 74.7 79.5 78.2
TRAN(2023)[31] S 94.3 90.7 93.6 77.7 85.0 81.6
DRNet(2023)[32] S 93.4 90.6 95.8 81.6 83.6 83.0
CASR-DRNet(2023)[58] S 95.1 92.3 95.3 83.0 85.0 89.2
I2C2W(2023)[33] S 94.3 91.7 95.0 82.8 83.7 93.1
Xia et. al.(2023)[34] S 91.9 91.3 96.1 94.2 80.6 78.0 83.4 88.8
HVSI(2024)[35] S 96.7 94.9 97.4 86.2 89.2 95.5
Context-relevant
SEED(2020)[14] S 93.8 89.6 92.8 80.0 81.4 83.6
VisionLAN(2021)[8] S 95.8 91.7 95.7 83.7 86.0 88.5
ABINet(2021)[9] S 96.2 93.5 97.4 86.0 89.3 89.2
ABINet(2021)[9] R 98.6 98.2 98.0 97.8 90.5 88.7 94.1 97.2
ConCLR(2022)[36] S 96.5 94.3 97.7 85.4 89.3 91.3
MATRN(2022)[10] S 96.6 95.0 97.9 95.8 86.6 82.8 90.6 93.5
PARSeq(2022)[12] S 97.0 93.6 97.0 96.2 86.5 82.9 88.9 92.2
PARSeq(2022)[12] R 99.1 97.9 98.3 98.4 90.7 89.6 95.7 98.3
STRT(2023)[11] S 97.6 95.7 97.6 86.7 90.1 94.9
HAAP S 97.7 95.1 97.2 96.4 87.8 85.6 90.7 95.1
HAAP R 99.3 98.2 98.7 98.8 91.3 90.7 97.4 99.3
IV-B Implementation Details
The entire training and testing is implemented in two RTX 3090 GPUs with 24GB of RAM based on the Pytorch library [59]. The training is performed with a batch size of 1024 for 63,630 iterations i.e. 20 epochs on a real dataset of 3,257,585 samples and 4 epochs on a synthetic dataset of 16.89M samples. Following previous work [12], the Adam [60] optimizer is used along with the 1cycle [61] learning rate scheduler with an initial learning rate of $57e-$ 5. We set a maximum label length of $T=25$ and used a character set of size $S=94$ containing mixed-case alphanumeric characters and punctuation. The image is enhanced, resized, and normalized. First, the image enhancement is done using a three-layer RandAugment [62] operation where sharpening is tuned to Gaussian Blur and Poisson Noise. Second, the image is unconditionally resized to $128\times 32$ pixels, and patches of $8\times 4$ size are used. For the model testing phase, we use a 36-character set i.e. containing numbers and 26 letters. We use word accuracy as the main evaluation metric i.e., the prediction is considered correct if and only if the characters match at all positions.
Remark The highlighted red and red lines in the visualization results indicate error recognition and omissions, respectively. The best performance result is shown in bold font.
IV-C Comparison with State-of-the-Arts
We compare HAAP with recent methods on 6 public benchmarks, and the quantitative results are listed in Table. II. We can see that HAAP outperforms the benchmark Parseq [12], realizing SOTA performance. The recognition performance in IC15 (incident scene text), SVTP (perspective scene text), and CUTE80 (curved text line image) are improved by $1.1\%$, $1.7\%$, and $1\%$ respectively. Our model achieves $99.3\%$ recognition accuracy on the IIIT5K (distorted, low-resolution scene text) dataset. This is an improvement of $2.6\%$ compared to the latest method, HVSI [35]. This verifies the excellent performance of HAAP on irregular text datasets, which contain a large number of low-quality images such as noisy and blurred images.
Besides validating the model on small-scale public benchmarks, we evaluate HAAP on three larger and more challenging recent benchmarks consisting of irregular text with different shapes, low-resolution images, rotations, and occlusions. The results are shown in Table. III. It achieves $80.6\%$, $85.5\%$, and $85.8\%$ outperforms the previous methods in terms of accuracy, respectively. Representative visible results are shown in Fig. 6, which demonstrates that our method is sufficiently robust to occlusion and text direction variability. In detail, we find that HAAP and recent methods are effective in solving complex background problems caused by perspective shifts or uneven lighting. However, artistic lettering and occlusion can lead to errors in previous work. The independence of heterogeneous data leads to ambiguous judgments by the recognizer, which can be effectively avoided by HAAP. Furthermore, HAAP can protect against additional distortions and ambiguities. For example, although the second half of ”CORNER” in Fig. 6 is ambiguous, our model can still reason effectively. HAAP also recognizes that the ”R” in ”SHARP” in Fig. 6 relies on AR reasoning with adaptive permutations.
TABLE III: The comparison of word accuracy on three challenging datasets. “S” and “R” denote Synthetic and real data, respectively.
Method Train data COCO ArT Uber
CRNN(2016)[22] S 49.3 57.3 33.1
TRBA(2021)[30] S 61.4 68.2 38.0
ViTSTR(2021)[29] S 56.4 66.1 37.6
ABINet(2021)[9] S 57.1 65.4 34.9
PARSeq(2022)[12] S 64.0 70.7 42.0
HAAP S 65.0 71.7 43.2
CRNN(2016)[22] R 62.2 66.8 51.0
TRBA(2021)[30] R 77.5 82.5 81.2
ViTSTR(2021)[29] R 73.6 81.0 78.2
ABINet(2021)[9] R 76.5 81.2 71.2
PARSeq(2022)[12] R 79.8 84.5 84.1
HAAP R 80.6 85.5 85.8
Figure 5: Visual comparison with other methods in terms of complexity and efficiency.
In addition, Fig. 5 and Table. IV shows the trade-off between accuracy and cost (parameters, FLOPS, and latency). For concise representation, we use $Avg$ to denote the average accuracy, which is computed as a weighted average based on the number of datasets. HAAP achieves the highest average word accuracy and exhibits competitive cost quality in all three metrics. Compared to PARSeq, HAAP uses far fewer parameters and FLOPs. In terms of latency, HAAP achieves $Avg-6$ of 95.0% with 15.2 ms/image, which outperforms previous methods.
TABLE IV: Comparison with other methods in terms of complexity and efficiency. $AVG-6$ represents the average accuracy in the six benchmarks.
Method Parameters(M) Flops(G) Latency(ms) Avg-6(%)
TRBA[30] 49.6 10.9 13.0 84.0
ABINet[9] 36.9 7.3 33.9 93.7
ViTSTR[29] 85.8 17.6 9.8 85.2
XIA et. al.[34] 87.2 12.6 37.6 88.4
PARSeq[12] 23.8 6.3 17.5 94.0
STRT[11] 57.9 20.6 31.8 92.8
Ours 23.8 3.2 15.2 95.0
Figure 6: Comparison of qualitative results on challenging textual data including blurring, distortion, occlusion, low resolution, perspective-shifting, and multi-directionality.
IV-D Ablation study
To evaluate the effectiveness of the proposed HAAP with IPN and CHA, we conduct ablation studies on nine benchmark datasets. In these experiments, all models are trained using real datasets with consistent settings of their model hyper-parameters.
The effectiveness of IPN To study the impact of adaptive permutations generated by IPN, we compare the sequence modeling strategies using left-to-right, bi-directional, and PLM. From Table. V, we find that the model using the bidirectional can achieve about $3.4\%$ improvement in $Avg-9$ compared to the left-to-right strategy. This proves that the double amount of information brought by bidirectional sequence modeling can effectively improve the correlation ability between sequence contexts. The introduction of PLM allows the model to achieve $90.12\%$ performance in terms of $Avg-9$ along with an additional 4 sets of random permutations. Immediately, the model with IPN achieves $0.28\%$ performance gain in terms of $Avg-9$ compared to the model with PLM. Meanwhile, as shown in Fig. 7, the model with IPN outperforms the introduction of PLM in terms of stability. This indicates that adaptive multi-directional sequence modeling can avoid the effects caused by the grammatical structure of words or by external disturbances. It also prevents uncertainty with human interference. The Comparison of qualitative results is shown in Fig. 8.
TABLE V: Ablation study of the IPN.$Avg-9$ represents the average accuracy in nine datasets. $P_{left2right}$ and $P_{right2left}$ represent the left-to-right and right-to-left permutation.
$P_{left2right}$ $P_{right2left}$ PLM IPN Avg-9(%)
- - - 86.63
- - 89.99
- - - 90.12
- - - 90.40
Figure 7: Comparison of Loss (Left) and Valuation (Right) curves with IPN or PLM.
The effectiveness of CHA Considering that the unified gradient flow allows the modules to interact with each other, we evaluate CHA and MHA while evaluating their impact in different sequencing sequences. As shown in Table. VI, we find that the model with joint PLM and MHA (PLM+MHA) requires additional IR to improve the generalization. The PLM+CHA model achieves $90.15\%$ performance in terms of $Avg$ without additional IR, outperforming the performance of the PLM+MHA model with 3 number of refinement iterations ($90.12\%$). In addition, compared to the PLM+MHA model, the IPN+CHA model achieves a performance gain of $0.59\%$ without the need for refinement iterations. Besides, we find that the contribution of IR is minimal by introducing CHA in any sequence modeling (PLM or IPN). Because of the uniform gradient flow, the IPN+CHA model achieves the performance of $90.40\%$ and gets a performance gain of $0.25\%$ compared to the PLM+CHA model. This suggests that CHA can enhance the model generalization by providing effective positive feedback to the adaptive permutation.
Figure 8: The qualitative results of the IPN.
TABLE VI: The ablation study of CHA.
PLM IPN MHA CHA number of IR (Avg-9.(%))
0 1 2
- - 89.81 89.96 90.12
- - 90.15 90.17 90.15
- - 89.97 90.21 90.28
- - 90.40 90.43 90.42
The qualitative results of IPN and CHA are shown in Fig. 9. We find that the introduction of IPN and CHA makes the model not only effective for arbitrary orientation character recognition but also has strong generalization ability for low-quality and occlusion.
Figure 9: Comparison of qualitative results.
IV-E Limitation and discussion
Although the proposed method can accurately recognize complex texts of arbitrary shapes in most cases, it still struggles with four kinds of tiny texts (image pixels not exceeding 2000): i:) front and back view severely confused; ii:) front and back text sticking together caused by blurring over long distances; iii:) shape-shifting with multiple occlusions; and iv:) alternating overlapping. Fig. 10 shows examples of such failure cases. The model loses attention on details during learning as the representation of character features declines in sequences.
Figure 10: Visualization results of some failure cases.
V Conclusion
To improve the autoregressive generalization based on internal LM, this paper proposes HAAP to enhance the location-context-image interaction capability by designing IPN and CHA. First, IPN is used to parse sequence meaning and syntactic structure by dynamically specifying token dependencies with an adaptive attention mask. The adaptive AR subset increases the diversity of the training data and prevents model dependency on specific sequences. The training overhead of PLM can be reduced and training fitting oscillations is avoided. Second, CHA module for hierarchical feature processing is proposed to exploit position-context-image semantic dependencies in the sequence without IR. The experimental results validate the effectiveness and the SOTA performance of of HAAP in terms of accuracy, complexity, and latency. In the future, we expect to extend HAAP to end-to-end detection and recognition, where the detector can be embedded in the visual decoder since the strong region perception is exhibited using the self-attention mechanism.
Acknowledgment
This work was supported by the Natural Science Fund of China under Grants 62171135, Industry-University Research Project of Education Department 2022.

### v2

@@
 Abstract
-Internal Language Model (LM)-based methods use permutation language modeling (PLM) to solve the error correction caused by conditional independence in external LM-based methods.
-However, random permutations of human interference cause fit oscillations in the model training, and Iterative Refinement (IR) operation to improve multimodal information decoupling also introduces additional overhead.
-To address these issues, this paper proposes the Hierarchical Attention autoregressive Model with Adaptive Permutation (HAAP) to enhance the location-context-image interaction capability, improving autoregressive generalization with internal LM.
-First, we propose Implicit Permutation Neurons (IPN) to generate adaptive attention masks to dynamically exploit token dependencies.
-The adaptive masks increase the diversity of training data and prevent model dependency on a specific order.
-It reduces the training overhead of PLM while avoiding training fit oscillations.
-Second, we develop Cross-modal Hierarchical Attention mechanism (CHA) to couple context and image features.
-This processing establishes rich positional semantic dependencies between context and image while avoiding IR.
-Extensive experimental results show the proposed HAAP achieves state-of-the-art (SOTA) performance in terms of accuracy, complexity, and latency on several datasets.
+Scene Text Recognition (STR) is challenging in extracting effective character representations from visual data when text is unreadable.
+Permutation language modeling (PLM) is introduced to refine character predictions by jointly capturing contextual and visual information.
+However, in PLM, the use of random permutations causes training fit oscillation, and the iterative refinement (IR) operation also introduces additional overhead.
+To address these issues, this paper proposes the Hierarchical Attention autoregressive Model with Adaptive Permutation (HAAP) to enhance position-context-image interaction capability, improving autoregressive LM generalization.
+First, we propose Implicit Permutation Neurons (IPN) to generate adaptive attention masks that dynamically exploit token dependencies, enhancing the correlation between visual information and context.
+Adaptive correlation representation helps the model avoid training fit oscillation.
+Second, the Cross-modal Hierarchical Attention mechanism (CHA) is introduced to capture the dependencies among position queries, contextual semantics and visual information.
+CHA enables position tokens to aggregate global semantic information, avoiding the need for IR.
+Extensive experimental results show that the proposed HAAP achieves state-of-the-art (SOTA) performance in terms of accuracy, complexity, and latency on several datasets.
 Index Terms:
@@
 I Introduction
-Scene Text Recognition (STR) aims to transcribe the text in an image into a computer-readable text format i.e., to recognize the localized text regions.
+Scene Text Recognition (STR) aims to transcribe text in an image to a computer-readable text format, i.e., to recognize the localized text regions.
 Scene text with rich information plays a vital role in a range of applications such as visual quizzing, autonomous driving, image retrieval, augmented reality, retail, education, and visually impaired devices [40, 42, 26, 5, 23].
@@
 The dashed line represents the interpretation of the mask generation: the visual information guides the adaptive mask to learn the inter-correlation of the contextual positions.
-In contrast to Optical Character Recognition (OCR) in documents with more homogeneous text attributes, STR has to deal with irregular text, such as rotation, bending, blurring, or occlusion.
-In addition, noise and distortion issues make STR challenging.
-STR is primarily a visual task, involving two different modalities: image and text.
-Early work still tends to rely on a backbone pre-trained on unimodal data.
-Learning from data [6, 7, 8, 9, 10, 11, 12, 13] is achieved by combining visual features with semantic priors from word representation models [14] or dictionaries [15] as well as sequence modeling [16].
-However, there are inherent limitations in models based on visual information.
-Visual features alone are not sufficient in cases where the text part is unreadable, e.g., occlusion.
-Additionally, STR is a One-shot task that has data independence.
-And, textual semantics can assist visual models in recognizing heterogeneous data.
-Zhao et al.
-verified that language modeling (LM) can effectively focus on text and accurately determine the meaning of words [13].
-LM demonstrates the special ability to perceive and understand text in natural images.
-For that, the combination of LM and visual models for sequence modeling is the current mainstream [8, 9, 10, 11, 12, 13].
-It is divided into two categories: external LM-based and internal LM-based.
-External LM-based schemes essentially accomplish STR by introducing external LMs to assist in correcting visual model reasoning [8, 9, 10, 11].
-External LM causes error correction own to the conditional independence of the input.
-The condition assumes that each pixel or character is independent, regardless of their interactions.
-However, the reality is that the order and spatial position of the characters about each other affect their meaning and grammatical structure.
-Algorithms that only consider each character by itself and ignore interactions can lead to prediction errors, especially when the input text has a complex structure and arrangement.
-Therefore, Permutation Language Modeling (PLM) is employed to improve the generalization of autoregressive (AR) modeling by parsing sequence meaning and syntactic structure [12, 13].
-PLM models can be viewed as a collection of AR models with shared architecture and weights.
-Token dependencies are specified dynamically through the use of an attention mask.
-Although randomized permutation increases the diversity of data, this strategy is accompanied by training fit oscillations since the random subset has the uncertainty of human interference.
-We also find that random permutation leads to unstable convergence of the model during training.
-In addition, if the model only performs well for specific alignments, it leads to a degradation of the generalization performance.
-Next, recent approaches have used Transformer encoder-decoder based for cross-modal visual and semantic information fusion [8, 9, 10, 11, 12, 13].
-However, They need to optimize the predictions through additional Iterative Refinement (IR).
-Essentially, internal LM is introduced in the decoder can eliminate IR, which is only an additional decoding because the visual and LM are jointly parsed.
-Considering that noisy inputs still have errors from the output of the visual model, IR is proposed to refine predictions from visual and language cues in external LM-based scenarios.
-Therefore, recent internal LM-based schemes do not propose an efficient way to couple cross-modal information in the decoder but rely on additional IR.
+In contrast to Optical Character Recognition (OCR) in documents with more homogeneous text attributes, STR poses significant challenges due to irregular text, including rotated, curved, blurred, and occluded characters.
+STR is primarily a vision-based task that involves two different modalities: image and text.
+Early works use a model pre-trained on image data and directly classify characters according to visual information of the text region without incorporating context, i.e., the meaning of text [36, 55].
+However, visual information alone is insufficient in cases where the text is unreadable, e.g., occluded, blurred, or with various fonts and spatial layouts [4].
+By language modeling (LM), text is introduced to aid recognition as text carries rich contextual information.
+Modeling the dependencies between the visual information and context and using autoregressive sequence modeling to predict character tokens [47, 11, 29, 49, 4, 59, 51].
+Current LM-based approaches can be divided into two categories: external LM-based [47, 11, 29, 49] and internal LM-based [4, 59, 51].
+The external LM-based approach introduces an independent LM to correct the token prediction of the visual model by explicitly modeling the context.
+In contrast, the internal LM-based approach embeds LM in the visual model to exploit joint visual information and context.
+External LM is used as a spell checker to rectify the visual prediction via unidirectional sequence modeling.
+The conditional independence of the LM from the input may cause it to correct the predictions erroneously [11, 49].
+Therefore, the internal LM-based approach applies Permutation Language Modeling (PLM) to enhance the correlation between the input image and context, improving the generalization of sequence modeling [4, 59].
+It can be seen as an ensemble of autoregressive models with shared architecture and weights [4].
+Using attention masks to dynamically specify token dependencies, PLM allows the decoder to perform sequence modeling of characters in arbitrary orders without relying on specific sequence order [59].
+Compared to unidirectional modeling, such as left-to-right (LTR) and right-to-left (RTL), the model can capture more information about token dependencies.
+However, there are still two problems: 1:) PLM produces a training fit oscillation problem due to the introduction of randomized permutation.
+Randomized permutation brings stochastic token dependencies that cause the model to have an unstable fitting process in decoding, which is defined as training fit oscillation in this paper.
+This makes decoding inefficient and requires multi-training to determine the best model.
+2:) Iterative refinement (IR) was initially proposed in external LM-based approaches [47, 11, 29, 49] and then in internal LM-based ones [4, 59, 51] to aggregate global information from visual and language models to refine prediction.
+However, internal LM-based models have adopted an attention-based encoder-decoder architecture that integrates visual and language modeling.
+It means that contextual semantics can be refined from visual information in the decoder instead of using IR, which is inefficient and increases decoding time and complexity.
 Figure 2: The basic flow of STR.
-(a) Visual feature coding and decoding (b) Visual-contextual representation based on external LM.
-(c) Visual-contextual representation based on internal LM.
-(d) Internal LM-based visual-contextual representation without Iterative Refinement (IR) (Ours).
-This paper proposes the Hierarchical Attention AR Model with Adaptive Permutation (HAAP) to address these issues.
-The motivation is depicted.
-First, the use of an adaptive permutation instead of a random permutation scheme increases the diversity of the training data and prevents dependence on a specific order.
-It reduces the training overhead of the PLM while avoiding training fitting oscillations (as shown in Fig.
+(a) Visual feature coding and decoding (b) Joint visual and context representation based on external LM.
+(c) Joint visual and context representation based on internal LM.
+(d) Internal LM-based visual-context representation without Iterative Refinement (IR) (Ours).
+To avoid the training fit oscillation and the use of IR, this paper proposes the Hierarchical Attention autoregressive model with Adaptive Permutation (HAAP) to improve the efficiency of cross-modal information decoding.
+In particular, first, we propose Implicit Permutation Neurons (IPN) to establish an adaptive attention mask in decoding by jointly considering the visual information and context (as illustrated in Fig.
 1).
-Second, Multi-head Attention (MHA) is introduced to hierarchically couple cross-modal information.
-Hierarchical processing improves the adaptation of visual features to semantics and thus enhances end-to-end joint parsing without additional IR (as shown in Fig.
+IPN adaptively exploits token dependencies, enhancing the input-output correlations within the model and preventing the training fit oscillation.
+Second, we develop the Cross-modal Hierarchical Attention mechanism (CHA) to integrate cross-modal information by establishing the dependencies among the position queries, context and the visual information using the adaptive attention mask provided by IPN.
+The contextual semantics are refined through a hidden layer, enabling each position token to be predicted based on global contextual semantics, rather than solely on preceding tokens.
+CHA significantly improves decoding efficiency and avoids the need for IR (as shown in Fig.
 2).
-First, inspired by traditional neurons and self-attentive mechanisms, this paper proposes an Implicit Permutation Neurons (IPN) to adapt the ordering between characters in the internal LM.
-The ordering and spatial location of the characters are considered rather than independently.
-IPN assigns a set of queries to the original left-to-right permutation to map it linearly into the high-dimensional space.
-Immediately, two sets of weight matrices are utilized for query weighting and inverse mapping.
-The process is still able to capture nonlinear relationships in the data with learnable weights.
-Second, this paper proposes Cross-modal Hierarchical Attention mechanism (CHA) to couple cross-modal information.
-Contexts embedded in the high dimensional space establish semantic dependencies through self-attention.
-The context interacts with the image features for the first time to context-image feature alignment.
-Next, the encoded location is matched with the context and then coupled with the image features again by mutual attention.
-Hierarchical feature processing drives the interactions of each item in the sequence while capturing the dependencies between locations using an image-guided context-assisted model.
-In addition, the location-context-image interaction is improved with location inter-correlation guided by adaptive mask and the uniform gradient flow.
-Therefore, the model avoids additional overhead by eliminating IR and achieves end-to-end parsing.
 The contributions of this paper can be summarized as follows:
-  • This paper proposes an internal LM-based HAAP to unify end-to-end coupling of location coding, image features, and context to improve the generalization ability of the model in various STR scenes.
-  • We propose a novel neuronal structure i.e.
-IPN to capture the dependencies between characters via adaptive masks efficiently.
-The model automatically learns the optimal alignment form by capturing the nonlinear relationships during the learning of weights without manual design.
-  • We propose hierarchical feature processing CHA to motivate the interaction of each item in the sequence, together with adaptive masks to improve the location-context-image interaction.
+  • This paper proposes an internal LM-based HAAP to end-to-end coupling of position, image, and context to improve the generalization ability of the model in various STR scenes.
+  • We propose a novel neuronal structure, i.e., IPN, to efficiently capture the token dependencies via an adaptive attention mask. The model automatically learns contextual semantic representations.
+  • A hierarchical feature processing CHA is introduced to establish the position-context-image interaction.
+Enhancing the correlation between visual information and context and expanding the semantic receptive field of each position token.
   • HAAP achieves state-of-the-art (SOTA) results on the STR benchmark for all character sets as well as on larger, more difficult real datasets that include occluded and arbitrarily oriented text (as shown in Table.
@@
 In this section, we review current deep learning-based STR methods.
-Recent studies are mainly categorized into two directions, context-independent and context-relevant, based on whether textual semantics are considered or not.
+Recent studies are mainly categorized as context-independent and context-relevant.
 We summarize and discuss the research in these two directions separately.
 II-A Context-independent STR
-Considering that STR is essentially a visual task, the researchers intuitively represent the inputs by constructing a visual model based on deep learning.
-This is the context-independent scheme i.e., category inference by summarizing visual features.
-Further, context-independent STR can be categorized into parallel-based reasoning and sequential reasoning.
-The parallel-based scheme only uses the visual features for prediction without considering the relationship between characters.
-Its main application is Fully Convolutional Network (FCN) [17] to segment characters at the pixel level.
+Since STR is a visual task, the researchers intuitively represent the input using a visual model and perform the class prediction from visual features.
+The parallel prediction-based methods directly use Fully Convolutional Network (FCN) [25] to segment characters pixel by pixel without considering the relationship between characters.
 Liao et al [27] recognized characters by grouping segmented pixels into text regions.
-Wan et al [19] proposed an additional sequential segmentation map that records characters in the correct order.
-Multi-category classification is unable to accurately construct complete phrases because the output characters are conditionally uncorrelated with each other.
-Immediately, the scheme of sequence inference using Recurrent Convolutional Networks (RNN) was proposed to capture the correlation of characters [20].
-The most typical is the Connectionist Temporal Classification (CTC) [21] based scheme.
-The RNN models the sequential sequence of features extracted by a convolutional neural network (CNN) and is trained end-to-end using CTC loss [22].
-The RNN receives the current input and the previous hidden state at each time step and outputs the current hidden state.
-This sequential modeling structure allows the RNN to naturally adapt to the characteristics of sequential data, to parse the syntactic and semantic structures in the text.
+Wan et al [45] identified characters in the correct order using an additional segmentation map.
+Multi-class prediction cannot accurately construct a complete text instance because the output characters are conditionally uncorrelated.
+Immediately, sequence modeling-based methods use Recurrent Convolutional Networks (RNN) to capture the correlation of characters [24].
+The most typical is the Connectionist Temporal Classification (CTC)-based scheme [12].
+The RNN models the feature sequence extracted by a convolutional neural network (CNN) and is end-to-end trained using CTC loss [35].
+The RNN receives the current input and the previous hidden state at each time step and outputs the current representation.
+This sequence modeling structure allows the RNN to naturally adapt to the characteristics of sequential data, to capture the linguistic rules and semantics in the text.
 RNN is also used in a series of Attention-based [3] schemes.
-Several attempts have explored the enhancement of image encoding and decoding by employing different attentional mechanisms [24, 25, 26].
+Several attempts have explored the enhancement of image encoding and decoding by employing different attention mechanisms [6, 19, 20].
 Character classification by aligning features and character positions, or converting STR into a multi-instance learning problem [55, 14].
-Sequence modeling was given a new lease of life with the introduction of transformer [16, 28].
+Sequence modeling-based methods have made great progress with the introduction of the Transformer [43, 10].
 Transformer was then proposed as an alternative to RNN for sequence modeling [1, 2, 48, 21, 52, 50, 9].
 However, context-independent methods that rely only on image features for prediction cannot address recognition in low-quality images, especially being less robust to corruption such as occlusion, distortion, or incomplete characters.
-This limitation motivates the use of language semantics to make recognition models more robust.
+This limitation motivates using context to make recognition models more robust.
 II-B Context-relevant STR
-Context-dependent STR is a typical cross-modal fusion scenario since the internal interactions between vision and language are constructed.
-Considering that textual semantics can assist visual models in processing heterogeneous data, the community has recently appeared with a large number of researches to assist recognition by constructing LMs [14, 8, 9, 12, 10, 36, 13, 11].
-VisionLAN [8] introduced character masking to guide the visual model to refer to language information in the visual context.
-MATRN [10] referenced context to enhance visual and semantic features based on spatial coding.
-STRT [11] designed an iterative text Transformer to predict the probability distribution in a sequence of characters.
-It is shown that robust models can be constructed by introducing external LMs for explicit modeling.
-Considering that character sequences are usually modeled in a left-to-right manner, the researcher develops the integrated model to capture twice the amount of information through the bi-directional LM [7, 37, 6, 38].
+Context-dependent STR is a typical cross-modal fusion process since the internal interactions between vision and context are constructed.
+Considering that contextual semantics can assist visual models in processing heterogeneous data, the community has recently appeared with a large number of research studies using LM to assist recognition [32, 47, 11, 4, 29, 57, 59, 49].
+VisionLAN [47] introduced character masking to guide the visual model to refer to text information in the visual context.
+MATRN [29] referenced context to enhance visual semantics based on spatial coding.
+STRT [49] designed an iterative text Transformer to predict the probability distribution in character sequences.
+It is shown that robust recognition models can be constructed by introducing external LMs for explicit modeling.
+Considering that character sequences are usually modeled in a left-to-right manner, the researcher develops the integrated model to capture twice the amount of information through the bi-directional LM [55, 36].
 SRN [55] combined the features of two unidirectional Transformers for prediction, which resulted in twice as expensive both computationally and parametrically.
 Therefore, ABINet [11] used a novel bidirectional completion network based on bidirectional feature representation.
-However, the conditional independence of external LM on image features limits its performance and makes error corrections.
-Subsequently, the internal LM was re-proposed for implicit language modeling to improve visual-contextual interaction.
-PARSeq [12] used PLM rather than standard AR modeling to learn internal LMs.
-It supports flexible decoding by using a parameterization that decouples the decoding location from the input context.
-CLIP4STR [13] introduced an encoder-decoder to process image and text information.
-The encoder is inherited from CLIP [39], while the decoder uses Transformer.
-Internal LR-based schemes mitigate the problem of noisy inputs as well as increase the diversity of the data by dynamically specifying the sequence order through the use of an attention mask.
-However, the strategy of using randomized permutations introduces training fitting oscillations due to the uncertainty of randomized permutations with human interference.
-Furthermore, most of these context-dependent STR models use Transformer-based encoder-decoders to implement with IR.
-However, internal LM-based schemes rely on IR to address noisy inputs in external LM-based models.
-Thus, the decoding strategy of LM model-based schemes at this stage is inefficient and introduces additional computational effort.
-This paper proposes to use an adaptive sequence ordering scheme to increase the diversity of the training data and prevent the dependence of the model on a specific order.
-To further enhance the efficiency of cross-modal information coupling at the decoder side, MHA is developed to Hierarchical coupling information, improving the semantic adaptation of visual features.
-Ultimately, end-to-end joint parsing is enhanced to avoid additional IR.
+STRT [49] used LM to capture the linguistic rule and refine the position token prediction.
+However, the conditional independence of external LM from the input image limits its performance and makes the model erroneously rectify the correct predictions.
+Subsequently, the internal LM was proposed for implicit language modeling to improve image-context interaction.
+PARSeq [4] used PLM to enhance the generalization of internal LM.
+It supports flexible decoding by dynamically specifying the dependence of position tokens and then decoding the position tokens from the input.
+CLIP4STR [59] introduced CLIP [33] and used PLM to construct an encoder-decoder framework to integrate image and text information.
+PLM-based methods mitigate the problem of noisy inputs and increase the diversity of the data by dynamically specifying the sequence order.
+However, the use of randomized permutations introduces the problem of training fit oscillation.
+Furthermore, most internal LM-based models employ an attention-based encoder-decoder architecture to integrate visual information with contextual semantics.
+Nevertheless, due to inefficient decoding, these models typically adopt IR to refine prediction, at the cost of increased computational complexity.
+Randomized token dependencies make an insufficient contextual semantic representation.
+Since each position token prediction only relies on the semantics preceding the tokens, the model needs to use IR further to aggregate global contextual semantics.
+This paper proposes to adopt an adaptive sequence modeling strategy that promotes training data diversity and effectively captures the interdependence between visual and contextual information.
+To further enhance the efficiency of cross-modal information decoding, the Multi-head attention mechanism (MHA) is used to hierarchically model relationships among position queries, contextual semantics, and visual information.
+The correlation between position and context is enhanced, enabling each token to access global contextual semantics and avoiding the need for IR.
 Figure 3: The pipeline of HAAP.
@@
 III Methodology
-In this section, we present in detail the principles and framework of our proposed Hierarchical Attention AR Model with Adaptive Permutation (HAAP), including Implicit Permutation Neurons (IPN) and Cross-modal Hierarchical Attention mechanism (CHA).
+In this section, we present in detail the principles and framework of our proposed Hierarchical Attention autoregressive Model with Adaptive Permutation (HAAP), including Implicit Permutation Neurons (IPN) and Cross-modal Hierarchical Attention mechanism (CHA).
 III-A Overview
@@
 3.
-In the encoding phase, the image and text inputs are represented as a series of patches and semantic tokens.
+In the encoding phase, the image and text inputs are represented as a series of patches and context tokens.
 Subsequently, MHA is used for hierarchical image-context interaction and decoding.
-Specifically, first, the Transformer encoder [28] is utilized to establish the internal feature associations of the image patches.
-Second, the IPN assigns bi-directional and adaptive permutation masks to the textual context for adjusting sequence order.
-The contexts embedded into the high-dimensional features are first established by self-attention to establish internal associations.
-Then MHA is used for context and visual information interaction and position query association.
-Third, the associated information further interacts with the visual information by utilizing mutual attention.
-Next, the position query is linearly decoded into contextual output.
-Figure 4: Illustration of a ViT layer from Dosovitskiy et al . [28]. $LN$ pertains to layer normalization. $MLP$ represents Multilayer Perceptron.
+Specifically, first, the Transformer encoder [10] is used to establish the internal feature representation and correlation of the image patches.
+Second, the IPN assigns bi-directional and adaptive attention masks to the context for adjusting the order of autoregressive sequence modeling.
+The first attention stage establishes the dependencies between the visual information and context with attention masks, producing the contextual semantics.
+The second attention stage encodes the contextual semantics and position through MHA into position queries.
+Finally, guide with visual information, position queries autoregressive decode the position token per iteration.
+Figure 4: Illustration of a ViT layer from Dosovitskiy et al. [10]. LNLN pertains to layer normalization. MLPMLP represents Multilayer Perceptron.
 III-B Encoder
-Vision encoder Considering that ViT [28] divides images into patches and uses full connectivity to learn global relationships can better extract global and complex visual information than traditional CNNs [40], HAAP follows the original ViT structure to build the visual encoder, which consists of $L=12$ layers of Transformer encoder.
-All layers share the same architecture as shown in Fig.
-4.
-The output of the last encoder is subjected to a Layer Normalization (LN).
-Formally, first, the image of size $H\times W\times C$ (height $H$, width $W$, channels $C$) is reshaped into a series of flat 2D patches $x_{p}\in\mathbb{R}^{N\times(P^{2}-C)}$, where $(P,P)$ is the resolution of each image block and $N=HW/P^{2}$ is the number of blocks generated.
+Vision encoder As shown in Fig.
+4, HAAP uses $L=12$ layers of the Transformer encoder [10] to build the visual encoder.
+The output of the last encoder is processed by layer normalization $LN$.
+First, the image of size $H\times W\times C$ (height $H$, width $W$, channels $C$) is reshaped into a series of flat 2D patches $x_{p}\in\mathbb{R}^{N\times(P^{2}-C)}$, where $(P,P)$ is the resolution of each image block and $N=HW/P^{2}$ is the number of blocks generated.
 The patch $x_{p}$ is flattened and mapped to $D$ dimensions using a trainable linear projection $E\in\mathbb{R}^{({P^{2}}C)\times D}$.
@@
 $z$ performs feature extraction in alternating MHA and Mult-Layer Perceptron (MLP).
-A normalized $LN$ is applied initially in each block to mitigate internal covariate bias during training.
+$LN$ is applied initially in each block to mitigate internal covariate shift during training.
 Additionally, residual concatenation is applied twice in the block to aid gradient propagation in the depth model, as:
@@
 where $MLP$ contains two nonlinear layers with GELU activation functions.
-The $z^{\prime}_{l}$ and $z_{l}$ are the visually encoded information.
-MHA is the extension of scaled dot-product attention to multiple representation subspaces or heads.
-In scaled dot-product attention, the similarity scores are computed using dot-product computation of the similarity between two $d_{k}$-dimensional vectors query $q$, and key $k$.
-Attention scores are transformed by the $d_{v}$-dimensional vector values $v$.
-The scaled dot product attention is defined as:
-$Attn(q,k,v)=softmax(\frac{qk^{T}}{\sqrt{d_{k}}})v,$ (4)
-further,
-$MHA(q,k,v)=Concat(head_{1},\cdots head_{h})W^{o},$ (5)
-where
-$head_{i}=Attn(qW^{q}_{i},kW^{k}_{i},vW^{v}_{i}),$ (6)
-where the computational cost of MHA is practically constant due to the dimension of the vector $d_{head}=d_{model}/h$, where $h$ is the number of heads.
-The $W^{q},W^{k},W^{v}\in\mathbb{R}^{d_{model}\times d_{head}}$ are obtained by connecting the heads and multiplying with the output projection matrix $W^{o}\in\mathbb{R}^{d_{model}\times d_{head}}$.
+The $z^{\prime}_{l}$ and $z_{l}$ are encoded visual information.
+MHA is the extension of scaled dot-product attention to multiple representation subspaces or heads [4].
 Context encoder The context uses a Tokenizer [4] for lowercase byte pair encoding BPE. The start, padding, and end of the text sequence are padded with [SOS], [PAD] and [EOS] tokens, respectively.
 III-C Implicit Permutation Neurons (IPN)
-Considering that STR traditionally relies on left-to-right $P_{left2right}$ or right-to-left $P_{right2left}$ unidirectional sequence modeling, we propose IPN to improve the correlation between inputs and outputs through multidirectional sequence modeling.
-Essentially, multidirectional sequence modeling considers internal correlations while considering bi-directionality to avoid the effects of word syntactic structures or external disturbances such as occlusions and distortions.
-For example, to predict the ”e” in the word ”clearance”, we consider ”cl_” and ”ecnara_”.
-” in the context of ”cl_ar”.
+Considering that STR traditionally relies on left-to-right $P_{left2right}$ or right-to-left $P_{right2left}$ unidirectional sequence modeling, we propose IPN to enhance the correlation between inputs and outputs through multidirectional sequence modeling.
+Essentially, multidirectional sequence modeling considers internal correlations of context while keeping attention to the bi-directional information flow to capture the word language rules and avoid the effort of visual interference (i.g, occlusion and distortion).
+For example, to predict the first “e” in the word “clearance,” we consider “cl_,” “ecnara_,” and the context of “cl_ar”.
 IPN provides an attention mask during the attention operation to generate dependencies between the input context and the output without actually replacing the text labels.
@@
 [E] 1 1 1 1
-Formally, first, given the $T$-length text labels $y=[y^{1},y^{2},\cdots,y^{T}]$, decompose the likelihood $[1,2,\cdots,T]$ according to the canonical order using the chain rule to build a left-to-right permutation of $P_{left2right}$ to obtain the model inference as:
+Formally, first, given the $T$-length text labels $y=[y^{1},y^{2},\cdots,y^{T}]$, decompose the likelihood $1,2,\cdots,T]$ according to the canonical order using the chain rule to build a left-to-right permutation of $P_{left2right}$ to obtain the model inference as:
 $\mathrm{log}p(y|x)=\sum_{t=1}^{T}{\mathrm{log}p_{\theta}(y_{t}|y_{<t},x)},$ (4)
@@
 where $M_{w}$ is the weighted transformation matrix.
-This process helps the model to parse the representation of the structure of the input data.
-Even though there is no explicit nonlinear activation function throughout the process, the model is still able to capture the nonlinear relationships in the data because of the learning of the weights.
-Fourth, the transformation matrix is inversely mapped to the sequence of scores $P_{score}$ via $P_{value}$, as:
+This process helps the model capture the input representation.
+Even though there is no explicit nonlinear activation function throughout the process, the model can still capture the nonlinear relationships in the data because of the learning of the weights.
+The transformation matrix is then mapped to the sequence of scores $P_{score}$ via $P_{value}$, as:
 $P_{score}=M_{w}\cdot P_{value}^{\intercal}.$ (7)
 Next, $P_{score}$ is normalized from largest to smallest to generate the actual ranking order.
-Immediately, we invert the left-to-right permutations $[1,2,.,T]$ and invert the adaptive permutation pairs to obtain twice the amount of information.
+Immediately, we invert the left-to-right $[1,2,.,T]$ and adaptive permutation to obtain twice the amount of information.
 III-D Cross-modal Hierarchical Attention mechanism (CHA)
-The standard Transformer parametric AR model is inefficient for decoding in multidirectional sequence modeling since the feature distribution predicted by the hidden state is independent of the target location [41].
+In multi-directional sequence modeling, the standard Transformer-based autoregressive model is inefficient since the feature distribution predicted by the hidden state is independent of the target position [54].
 Let us assume that we parameterize the next token distribution as:
@@
 where $h_{\theta}(x_{r<t})$ denotes the hidden representation of $x_{r<t}$ produced by the shared transformer network after proper masking.
-We use $rt$ and $r<t$ to denote the $t$-th and the first $t-1$ elements of the permutation set $r$.
-Since the same model parameter $\theta$ is shared between all decomposition sequences during training, $rt$ is expected to see that every possible element in the sequence is not equal to $x_{t}$.
-Note that the representation $h_{\theta}(x_{z<t})$ does not depend on the position it will predict, i.e.
-the value of $rt$.
-Two different target positions share the same model predictions.
-However, the ground truth distributions for the two positions are certainly different.
-To avoid this problem, we perform context and location encoding based on a two-stream attention mechanism to provide complete contextual information.
-We use two sets of hidden representations i.e., from content and query representations.
-Meanwhile, considering that context and location need to interact with visual features cross-modally for alignment, we adopt a hierarchical coupling approach, i.e., CHA, to consider the context-location-image association.
-In the first stage, context-visual attention is established after self-attentive encoding of sorted contexts, as:
-$Attn_{cv}=p+dropout[MHA(Attn_{c},z)],$ (12)
+We use $x_{rt}$ and $x_{r<t}$ to denote the $t$-th and the first t1t-1 elements of the permutation set $r$.
+Since the same model parameter $\theta$ is shared between all decomposition sequences during training, $x_{rt}$ is expected to see every possible element in the sequence.
+Note that the position token $x_{rt}$ to be predicted only depends on the representation $h_{\theta}(x_{r<t})$.
+If only using the initial contextual semantics learned from the multi-directional sequence modeling without considering the current input visual information, it is not possible to update the representation $h_{\theta}(x_{r<t})$ with global information, increasing the probability of error accumulation.
+For this, researchers employ IR to improve the representation of $h_{\theta}(x_{r<t})$ [11, 4].
+The IR uses the cloze mask to enhance the understanding of $h_{\theta}(x_{r<t})$ for the context because the cloze mask is able to provide the rest of the information other than the current position.
+To avoid this problem, we introduce CHA to hierarchically encode context and position by allowing context, position, and image to interact with each other through a two-stream attention mechanism.
+We use two sets of hidden representations, i.e., from contextual semantic and query representations.
+In the first stage, context-image attention is established, producing the context semantic representation, which encodes both the context and $x_{rt}$ itself tokens guided by visual information, as:
+$Attn_{cv}=Attn_{c}+dropout[MHA(Attn_{c},z_{l})],$ (9)
 where
 $Attn_{c}=c+dropout[MHA(c,c,M)],$ (10)
-where $c\in\mathbb{R}^{(T+1)\times d_{model}}$ is the contextual embedding with positional information and $M\in\mathbb{R}^{(T+1)\times(T+1)}$ is the attention mask.
-The total sequence length is increased to $T+1$ by using special separator markers.
-Next, MHA is used to establish positional-contextual-visual attention, as:
-$Attn_{f}=Attn_{p}+dropout[MHA(Attn_{p},z)],$ (14)
+where $c\in\mathbb{R}^{(T+1)\times d_{model}}$ is the context embeddings with position information and $M\in\mathbb{R}^{(T+1)\times(T+1)}$ is the attention mask.
+The total sequence length is increased to $T+1$ because of the use of special delimiter tokens ([B] or [E]).
+Next, MHA is used to establish position-context-image attention.
+as:
+$Attn_{f}=Attn_{p}+dropout[MHA(Attn_{p},z_{l})],$ (11)
 where
-$Attn_{p}=p+dropout[MHA(p,Attn_{c}v,M)].$ (15)
-The strategy of hierarchical feature processing prompts sequence interactions while utilizing image-guided context to assist the model in capturing dependencies between locations.
-The location-context-image interaction capability is boosted since adaptive permutation guides location interconnections.
-The robustness of the model is improved because the uniform gradient flow allows them to interact with each other during training.
+$Attn_{p}=p+dropout[MHA(p,Attn_{cv},M)],$ (12)
+where $p$ is the position queries.
+By updating the contextual semantic representation, each position token can gather the global information guided by the current input.
+CHA first helps the model capture dependencies between context and the visual information by allowing sequence interactions with the adaptive attention mask ($M$) provided by IPN.
+Then, the correlation between position and context is enhanced through context-image implicit guidance, thus getting rid of additional IR.
+The robustness of the model is improved because the uniform gradient flow allows IPN and CHA to interact with each other during training.
 Finally, the output logits $y^{\prime}$ are obtained as:
@@
 where $S$ is the size of the character set used for training.
-Additional characters are associated with the [E] tag (the end of the tagging sequence).
-The complete training loss for the optimization phase is the average of the $K$ cross-entropy loss $\mathcal{L}_{ce}$ with attention mask, as:
+Additional characters are associated with the [E] token (the end of the token sequence).
+In the optimization phase, the training loss is the average of the $K$ cross-entropy loss $\mathcal{L}_{ce}$ with attention mask, as:
 $\mathcal{L}=\frac{1}{k}\sum_{k=1}^{K}{\mathcal{L}_{ce}(y^{\prime}_{k},y)},$ (14)
@@
 We use IIIT 5k-word (IIIT5k) [28], CUTE80 (CUTE) [34], Street View Text (SVT) [46], SVT-Perspective (SVTP) [31], ICDAR 2013 (IC13) [16] and ICDAR 2015 (IC15) [15], COCO, Uber, ArT for testing.
-The dataset contains challenging texts such as multi-orientation, curvature, perspective blur, low resolution, distortion, and occlusion.
-Synthetic datasets MJ consists of 8.9 million photorealistic text images composed of artificial datasets through over 90k English dictionaries.
+The dataset contains challenging texts such as multi-orientation, curved, blurry, low resolution, distorted, and occluded text.
+Synthetic datasets MJ consists of 8.9 million photorealistic text images composed of synthetic datasets through over 90k English dictionaries.
 The composition of MJ consists of three parts background, foreground, and optional shadow/border respectively.
@@
 The text image is cropped from the natural image in which the synthesized text is embedded.
-Real-world datasets The COCO dataset has a total of 73,127 training and 9.8k test images containing non-text, legible text, illegible text, and occluded text images.
-Each image contains at least one instance of legible text.
+Real-world datasets The COCO dataset has a total of 73,127 training and 9.8k test images containing non-text, legible, illegible, and occluded text images.
+Each image contains at least one instance of readable text.
 RCTW contains 12,263 annotated images of the large-scale Chinese field dataset.
-The Uber dataset has a total of 127,920 training 60,000 test images collected from Bing Maps Streetside.
+The Uber dataset has 127,920 training 60,000 test images collected from Bing Maps Streetside.
 It contains house numbers and vertical and rotated text on signboards.
-The ArT dataset was created to recognize arbitrarily shaped text containing 32,028 training and 26,000 testing images of perspective, rotated, or curved arbitrarily shaped text.
-SVT is a large-scale street scene text dataset collected from streets in China totaling 41,439 text images of arbitrarily shaped natural scenes.
+The ArT dataset was created to recognize arbitrarily shaped text containing 32,028 training images and 26,000 test images of arbitrarily shaped, perspective, rotated, or curved text.
+SVT is a large-scale street scene text dataset collected from streets in China, including 41,439 text images of arbitrarily shaped natural scenes.
 MLT19 was created to recognize multilingual text.
@@
 The ReCTS dataset contains 26,432 irregular texts arranged in various layouts or written in unique fonts.
-TextOCR and OpenVINO are large datasets with very diverse images containing 818,087 and 2,071,541 images respectively.
+TextOCR and OpenVINO are large datasets with diverse images containing 818,087 and 2,071,541 images, respectively.
 They have complex scenes with multiple objects and text of different resolutions, orientations, and qualities.
-The IIIT5k dataset is a collection of 5000 natural scene and digitally born text images crawled from Google image search containing 2,000 images for evaluation and 3,000 images for testing.
+The IIIT5k dataset is a collection of 5000 natural scenes and digitally born text images crawled from Google image search, containing 2,000 images for evaluation and 3,000 images for testing.
 It has low resolution, multiple font styles, light and dark transformations, and projection distortion.
@@
 It contains 257 images for evaluation and 647 images for testing.
-Further, the SVTP dataset is a more complex collection of images containing numerous perspective texts totaling 645 images for testing.
+Further, the SVTP dataset is a more complex collection of images containing numerous perspective texts, totaling 645 images for testing.
 IC13 and IC15 were created for the ICDAR Robust Reading Contest.
@@
 IC15 contains 4,468 images for training and 2,077 images for evaluation.
-Many of the samples contain perspective, and blurred text.
+Many of the samples contain perspective and blurred text.
 TABLE II: The comparison of word accuracy on six common benchmarks in terms of both context relevant and independent. “S” and “R” denote Synthetic and real data, respectively.
@@
 HVSI(2024)[9] S 96.7 94.9 97.4 86.2 89.2 95.5
+CAM(2024)[53] S 97.4 96.1 97.2 87.8 90.6 92.4
 Context-relevant
@@
 STRT(2023)[49] S 97.6 95.7 97.6 86.7 90.1 94.9
+OTE(2024)[51] S 96.2 93.5 97.6 85.9 89.6 91.7
 HAAP S 97.7 95.1 97.2 96.4 87.8 85.6 90.7 95.1
@@
 IV-B Implementation Details
-The entire training and testing is implemented in two RTX 3090 GPUs with 24GB of RAM based on the Pytorch library [59].
-The training is performed with a batch size of 1024 for 63,630 iterations i.e.
+The entire training and testing is implemented in two RTX 3090 GPUs with 24GB of RAM.
+The training is performed with a batch size of 1024 for 31,815 iterations i.e.
 10 epochs on a real dataset of 3,257,585 samples and 4 epochs on a synthetic dataset of 16.89M samples.
-Following previous work [12], the Adam [60] optimizer is used along with the 1cycle [61] learning rate scheduler with an initial learning rate of $57e-$ 5.
+Adam [18] optimizer is used along with the 1cycle [39] learning rate scheduler with an initial learning rate of 7e57e-5.
 We set a maximum label length of $T=25$ and used a character set of size $S=94$ containing mixed-case alphanumeric characters and punctuation.
-The image is enhanced, resized, and normalized.
-First, the image enhancement is done using a three-layer RandAugment [62] operation where sharpening is tuned to Gaussian Blur and Poisson Noise.
-Second, the image is unconditionally resized to $128\times 32$ pixels, and patches of $8\times 4$ size are used.
+The image is augmented using a three-layer RandAugment [8] operation where sharpening is tuned to Gaussian Blur and Poisson Noise and resized to $128\times 32$ pixels.
+Each patch is set a size of $8\times 4$.
 For the model testing phase, we use a 36-character set i.e.
@@
 We use word accuracy as the main evaluation metric i.e., the prediction is considered correct if and only if the characters match at all positions.
-Remark The highlighted red and red lines in the visualization results indicate error recognition and omissions, respectively. The best performance result is shown in bold font.
+Remark The highlighted red and red lines in the visualization results indicate error prediction and lack of word, respectively. The best performance result is shown in bold font.
 IV-C Comparison with State-of-the-Arts
@@
 We can see that HAAP outperforms the benchmark Parseq [4], realizing SOTA performance.
-The recognition performance in IC15 (incident scene text), SVTP (perspective scene text), and CUTE80 (curved text line image) are improved by $1.1\%$, $1.7\%$, and $1\%$ respectively.
+The recognition performances in IC15 (incident scene text), SVTP (perspective scene text), and CUTE80 (curved text) are improved by $1.1\%$, $1.7\%$, and $1\%$ respectively.
 Our model achieves $99.3\%$ recognition accuracy on the IIIT5K (distorted, low-resolution scene text) dataset.
 This is an improvement of $2.6\%$ compared to the latest method, HVSI [9].
-This verifies the excellent performance of HAAP on irregular text datasets, which contain a large number of low-quality images such as noisy and blurred images.
-Besides validating the model on small-scale public benchmarks, we evaluate HAAP on three larger and more challenging recent benchmarks consisting of irregular text with different shapes, low-resolution images, rotations, and occlusions.
+This verifies the excellent performance of HAAP on irregular text datasets, which contain a large number of low-quality images such as distorted and blurred images.
+Besides validating the model on small-scale public benchmarks, we evaluate HAAP on three larger and more challenging recent benchmarks consisting of irregular text with different shapes, low-resolution, rotated, and occluded text.
 The results are shown in Table.
@@
 Representative visible results are shown in Fig.
-6, which demonstrates that our method is sufficiently robust to occlusion and text direction variability.
+5, which demonstrates that our method is sufficiently robust to occlusion and text direction variability.
 In detail, we find that HAAP and recent methods are effective in solving complex background problems caused by perspective shifts or uneven lighting.
@@
 5 relies on AR reasoning with adaptive permutations.
+Figure 5: Comparison of qualitative results on challenging textual data including blurring, distortion, occlusion, low resolution, perspective-shifting, and multi-directionality.
+Figure 6: Visual comparison with other methods in terms of complexity and efficiency.
 TABLE III: The comparison of word accuracy on three challenging datasets. “S” and “R” denote Synthetic and real data, respectively.
@@
 HAAP R 80.6 85.5 85.8
-Figure 5: Visual comparison with other methods in terms of complexity and efficiency.
-In addition, Fig.
-5 and Table.
-IV shows the trade-off between accuracy and cost (parameters, FLOPS, and latency).
-For concise representation, we use $Avg$ to denote the average accuracy, which is computed as a weighted average based on the number of datasets.
-HAAP achieves the highest average word accuracy and exhibits competitive cost quality in all three metrics.
-Compared to PARSeq, HAAP uses far fewer parameters and FLOPs.
-In terms of latency, HAAP achieves $Avg-6$ of 95.0% with 15.2 ms/image, which outperforms previous methods.
-TABLE IV: Comparison with other methods in terms of complexity and efficiency. $AVG-6$ represents the average accuracy in the six benchmarks.
+TABLE IV: Comparison with other methods in terms of complexity and efficiency. AVG6AVG-6 represents the average accuracy in the six benchmarks.
 Method Parameters(M) Flops(G) Latency(ms) Avg-6(%)
@@
 XIA et. al.[50] 87.2 12.6 37.6 88.4
-PARSeq[12] 23.8 6.3 17.5 94.0
+PARSeq[4] 23.8 4.3 17.5 94.0
 STRT[49] 57.9 20.6 31.8 92.8
-Ours 23.8 3.2 15.2 95.0
-Figure 6: Comparison of qualitative results on challenging textual data including blurring, distortion, occlusion, low resolution, perspective-shifting, and multi-directionality.
-IV-D Ablation study
-To evaluate the effectiveness of the proposed HAAP with IPN and CHA, we conduct ablation studies on nine benchmark datasets.
-In these experiments, all models are trained using real datasets with consistent settings of their model hyper-parameters.
-The effectiveness of IPN To study the impact of adaptive permutations generated by IPN, we compare the sequence modeling strategies using left-to-right, bi-directional, and PLM.
-From Table.
-V, we find that the model using the bidirectional can achieve about $3.4\%$ improvement in $Avg-9$ compared to the left-to-right strategy.
-This proves that the double amount of information brought by bidirectional sequence modeling can effectively improve the correlation ability between sequence contexts.
-The introduction of PLM allows the model to achieve $90.12\%$ performance in terms of $Avg-9$ along with an additional 4 sets of random permutations.
-Immediately, the model with IPN achieves $0.28\%$ performance gain in terms of $Avg-9$ compared to the model with PLM.
-Meanwhile, as shown in Fig.
-7, the model with IPN outperforms the introduction of PLM in terms of stability.
-This indicates that adaptive multi-directional sequence modeling can avoid the effects caused by the grammatical structure of words or by external disturbances.
-It also prevents uncertainty with human interference.
-The Comparison of qualitative results is shown in Fig.
-8.
-TABLE V: Ablation study of the IPN.$Avg-9$ represents the average accuracy in nine datasets. $P_{left2right}$ and $P_{right2left}$ represent the left-to-right and right-to-left permutation.
+Ours 20.9 3.6 15.2 95.0
+In addition, Fig.
+6 and Table.
+IV shows the trade-off between accuracy and cost (parameters, FLOPS, and latency).
+For concise representation, we use $Avg$ to denote the average accuracy, which is computed as a weighted average based on the number of datasets.
+HAAP achieves the highest average word accuracy and exhibits competitive cost-quality in all three metrics.
+Compared to PARSeq, HAAP uses far fewer parameters and FLOPs.
+In terms of latency, HAAP achieves Avg6Avg-6 of 95.0% with 15.2 ms/image, which outperforms previous methods.
+TABLE V: Ablation study of the IPN.Avg9Avg-9 represents the average accuracy in nine datasets. $P_{left2right}$ and $P_{right2left}$ represent the left-to-right and right-to-left permutation.
 $P_{left2right}$ $P_{right2left}$ PLM IPN Avg-9(%)
@@
 - - - 90.40
-Figure 7: Comparison of Loss (Left) and Valuation (Right) curves with IPN or PLM.
-The effectiveness of CHA Considering that the unified gradient flow allows the modules to interact with each other, we evaluate CHA and MHA while evaluating their impact in different sequencing sequences.
+IV-D Ablation study
+To evaluate the effectiveness of the proposed HAAP with IPN and CHA, we conduct ablation studies on nine benchmark datasets.
+In these experiments, all models are trained using real datasets with consistent settings of their model hyperparameters.
+The effectiveness of IPN We compare IPN with different sequence modeling principles (i.g, left-to-right, bi-directional, and PLM) to evaluate the effectiveness of adaptive sequence modeling.
+Table.
+V shows that the bidirectional model can achieve a performance gain of about $3.4\%$ on Avg9Avg-9 compared to the left-to-right model.
+Compared to the PLM model ($90.12\%$), the IPN model achieves a performance gain of $0.28\%$ in terms of Avg9Avg-9.
+To demonstrate the effectiveness of IPN in model training, we show the training and validation curves in five continuous training sessions in Fig.
+7.
+Meanwhile, we use early stopping to determine the iteration step of model convergence by validating that the increase of Avg9Avg-9 is less than $0.3\%$ in 2000 steps.
+The result is depicted in Tab.
+VI.
+Specifically, as shown in Fig.
+7, the IPN model can achieve more stable training and convergence than the PLM model in multi-training.
+As depicted in Tab.
+VI, the Avg9Avg-9 standard deviation (Std Dev) of the IPN model is $0.33$, which is much less than the PLM model ($1.45$).
+With IPN, adaptive sequence modeling can help the model avoid train fit oscillation and achieve convergence advantage with less training data and time costs than PLM.
+The qualitative results in Fig.
+8 show that the IPN model can effectively predict some complex cases influenced by background.
+(a) the model with PLM
+(b) the model with IPN
+Figure 7: Comparison of Loss (Left) and Valuation (Right) curves in five continuous training sessions.
+TABLE VI: Training schedule comparison in five continuous training sessions. Std Dev represents standard deviation.
+Training session Step of convergence Avg-9 (%) Sample count (M) Training hours
+PLM
+Session 1 15318 90.12 15.69 5.23
+Session 2 19591 88.89 20.06 5.79
+Session 3 18500 86.83 18.94 5.64
+Session 4 20091 87.62 20.57 5.86
+Session 5 999 0 10.23 2.6
+Mean 18375 88.37 18.82 5.63
+Std Dev 2143.53 1.45 2.19 0.28
+IPN
+Session 1 11636 90.37 11.92 4.67
+Session 2 12545 89.67 12.85 4.79
+Session 3 12545 89.76 12.85 4.79
+Session 4 13727 89.78 14.06 4.91
+Session 5 12545 90.31 12.85 4.79
+Mean 12599 89.98 12.91 4.79
+Std Dev 743.05 0.33 0.76 0.08
+Figure 8: The qualitative results of the IPN.
+The effectiveness of CHA To evaluate the effectiveness of CHA, we compare PLM-based and IPN-based models with CHA in different numbers of IRs.
 As shown in Table.
-VI, we find that the model with joint PLM and MHA (PLM+MHA) requires additional IR to improve the generalization.
-The PLM+CHA model achieves $90.15\%$ performance in terms of $Avg$ without additional IR, outperforming the performance of the PLM+MHA model with 3 number of refinement iterations ($90.12\%$).
-In addition, compared to the PLM+MHA model, the IPN+CHA model achieves a performance gain of $0.59\%$ without the need for refinement iterations.
-Besides, we find that the contribution of IR is minimal by introducing CHA in any sequence modeling (PLM or IPN).
-Because of the uniform gradient flow, the IPN+CHA model achieves the performance of $90.40\%$ and gets a performance gain of $0.25\%$ compared to the PLM+CHA model.
-This suggests that CHA can enhance the model generalization by providing effective positive feedback to the adaptive permutation.
-Figure 8: The qualitative results of the IPN.
-TABLE VI: The ablation study of CHA.
-PLM IPN MHA CHA number of IR (Avg-9.(%))
-0 1 2
-- - 89.81 89.96 90.12
-- - 90.15 90.17 90.15
-- - 89.97 90.21 90.28
-- - 90.40 90.43 90.42
-The qualitative results of IPN and CHA are shown in Fig.
-9.
-We find that the introduction of IPN and CHA makes the model not only effective for arbitrary orientation character recognition but also has strong generalization ability for low-quality and occlusion.
+VII, we find that both the PLM model and IPN model without CHA require additional IR to improve the autoregressive generalization.
+By using CHA, the PLM model achieves the Avg9Avg-9 of $90.15\%$ without IR, outperforming the one after iterating IR three times without CHA ($90.14\%$).
+Based on CHA, the IPN model achieves a performance gain of $0.43\%$ without IR, indicating the effectiveness of CHA in decoding position tokens.
+In addition, Fig.
+9 shows the trade-off between the Avg9Avg-9, the model computational complexity (FLOPs), and single-image latency (Latency) with the number of IRs.
+The model without CHA requires two times of IR to achieve the stable performance, while the one with CHA maintains the performance advantage regardless of IR.
+Thus, the use of CHA avoids the additional computational complexity and latency induced by IR.
+As shown in Fig.
+10, although IR enhances the confidence of the IPN model, it results in wrong predictions.
+In contrast, the IPN model with CHA can generate high-confidence and accurate predictions without IR.
+The results clearly show that CHA can improve the performance of the autoregressive model through position-context-image interactions without the need for additional contextual semantic aggregation through IR.
+The qualitative results in Fig.
+11 show that the CHA-based model effectively recognizes low-resolution and blurred texts.
+TABLE VII: The ablation study of CHA.
+Algorithm and complexity Number of IR (Avg-9.(%))
+0 1 2 3 4
+PLM w/o CHA 89.81 89.96 90.12 90.14 90.12
+IPN w/o CHA 89.97 90.21 90.28 90.30 90.31
+FLOPs(G) 3.38 3.54 3.70 3.86 4.02
+PLM with CHA 90.15 90.17 90.15 90.16 90.13
+IPN with CHA 90.40 90.43 90.42 90.42 90.41
+FLOPs(G) 3.62 3.81 4.00 4.19 4.38
+Figure 9: Average accuracy (Avg-9) in nine datasets and FLOPs of the model (Left) and single-image latency (Right) for each decoding scheme. The number of IR used is indicated for each point.
+Figure 10: Comparison of confidence heat maps.
+Red color indicates the highest score, and blue color indicates the lowest score.
+For each approach from top to bottom are the score region visualization, confidence level, and prediction results.
+Ground truths are shown below the original images.
 Figure 11: Comparison of qualitative results.
 IV-E Limitation and discussion
-Although the proposed method can accurately recognize complex texts of arbitrary shapes in most cases, it still struggles with four kinds of tiny texts (image pixels not exceeding 2000): i:) front and back view severely confused; ii:) front and back text sticking together caused by blurring over long distances; iii:) shape-shifting with multiple occlusions; and iv:) alternating overlapping.
+Although the proposed method can accurately recognize arbitrary-shaped texts in most cases, it still struggles with four kinds of tiny texts (the total number of pixels does not exceed 2000) due to the lack of visual representation of tiny text.: i:) target confused with the background; ii:) blurred long-text; iii:) deformed text under multi-occlusion; and iv:) overlapping text.
 Fig.
 12 shows examples of such failure cases.
-The model loses attention on details during learning as the representation of character features declines in sequences.
 Figure 12: Visualization results of some failure cases.
 V Conclusion
-To improve the autoregressive generalization based on internal LM, this paper proposes HAAP to enhance the location-context-image interaction capability by designing IPN and CHA.
-First, IPN is used to parse sequence meaning and syntactic structure by dynamically specifying token dependencies with an adaptive attention mask.
-The adaptive AR subset increases the diversity of the training data and prevents model dependency on specific sequences.
-The training overhead of PLM can be reduced and training fitting oscillations is avoided.
+This paper proposes HAAP to enhance the position-context-image interaction capability using IPN and CHA to improve autoregressive generalization.
+First, IPN captures the correlation between context and visual information by dynamically specifying token dependencies with an adaptive attention mask.
+Adaptive sequence modeling increases the diversity of the training data and prevents the model from training fit oscillation.
 Second, CHA module for hierarchical feature processing is proposed to exploit position-context-image semantic dependencies in the sequence without IR.
-The experimental results validate the effectiveness and the SOTA performance of of HAAP in terms of accuracy, complexity, and latency.
-In the future, we expect to extend HAAP to end-to-end detection and recognition, where the detector can be embedded in the visual decoder since the strong region perception is exhibited using the self-attention mechanism.
+The experimental results validate the effectiveness and SOTA performance of HAAP in terms of accuracy, complexity, and latency.
+In the future, we expect to extend HAAP to end-to-end detection and recognition, where the detector can be embedded in the visual decoder since position awareness is exhibited using the attention mechanism.
 Acknowledgment
-This work was supported by the Natural Science Fund of China under Grants 62171135, Industry-University Research Project of Education Department 2022.
+This work was supported by the NSFC (No. 62171135), Fujian Key Project (No.2023XQ004), Fujian Eagle-Scholar and Distinguished Project (No. 2022J06010), Fujian Natural Science Fund (No. 2024H6013).